OpenStudy (anonymous):

Please help, I keep messing up near the end: Find the point on the line -6 x + 5 y - 6 =0 which is closest to the point ( 4, -3 ).

OpenStudy (cwrw238):

slope of -6 x + 5 y - 6 =0.............(a) is 6/5 slope of a line perpendicular to the above is -5/6. equn. of line with this slope passing thru (4,-3): y --3 = (-5/6)(x - 4) y + 3 = (-5/6)(x-4)...................(b) solve (a) and (b) simultaneously will give you the required point.

OpenStudy (anonymous):

Thank you for your help, I'm sorry for the trouble I was trying another method, what do you mean by solving them simultaneously

OpenStudy (cwrw238):

find values of x and y using elimination or substitution b = 6y + 18 = -5x + 20 0r 5x + 6y - 2 = 0...........b -6 x + 5 y - 6 =0.............(a) multiply b by 6 and a by 5 and add to eliminate x

OpenStudy (anonymous):

perfect, thank you very much!

OpenStudy (cwrw238):

yw